# On the Subpacketization Level of the Banawan–Ulukus Multi-Message PIR Scheme

### v1

@@ II Proof of Theorem 1
For each $n\in[0:T(D-1)]$, define $c_{n}$ as in (5), so that
$(1+x+\cdots+x^{D-1})^{T}=\sum_{n=0}^{T(D-1)}c_{n}\,x^{n}.$ (20)
In particular, $c_{n}\geq 0$, since $c_{n}$ equals the number of $T$-tuples $(n_{1},\dots,n_{T})\in[0:D-1]^{T}$ such that $\sum_{l=1}^{T}n_{l}=n$.
@@
$\left(\sum_{t=0}^{D-1}\omega_{m}^{t}s^{-t}\right)^{T}=\sum_{n=0}^{T(D-1)}c_{n}\,\omega_{m}^{n}\,s^{-n}.$ (22)
@@
$\frac{(1-N^{-1})^{T}}{(1-\omega_{m}s^{-1})^{T}}=\sum_{n=0}^{T(D-1)}c_{n}\,\omega_{m}^{n}\,s^{-n}.$ (23)
@@
$\frac{1}{D}\sum_{m=0}^{D-1}\frac{(1-N^{-1})^{T}}{(1-\omega_{m}s^{-1})^{T}}$ $=\sum_{n=0}^{T(D-1)}c_{n}\,s^{-n}\left(\frac{1}{D}\sum_{m=0}^{D-1}\omega_{m}^{n}\right)$
$=\sum_{k=0}^{S}c_{kD}\,s^{-kD}$
$=\sum_{k=0}^{S}c_{kD}\,N^{-k},$ (24)
@@
Since (6) is a finite sum of nonnegative integer powers of NN with coefficients $c_{kD}\geq 0$ for all $k\in[0:S]$, it follows that $L$ is a polynomial in NN with nonnegative coefficients. Moreover, the highest power of NN occurs at $k=0$, and
$c_{0}=[x^{0}](1+x+\cdots+x^{D-1})^{T}=1,$

### v2

@@ II Proof of Theorem 1
For each $n\in[0:T(D-1)]$, define $a_{n}$ as in (5), so that
$(1+x+\cdots+x^{D-1})^{T}=\sum_{n=0}^{T(D-1)}a_{n}\,x^{n}.$ (20)
In particular, $a_{n}\geq 0$, since $a_{n}$ equals the number of $T$-tuples $(n_{1},\dots,n_{T})\in[0:D-1]^{T}$ such that $\sum_{l=1}^{T}n_{l}=n$.
@@
$\left(\sum_{t=0}^{D-1}\omega_{m}^{t}s^{-t}\right)^{T}=\sum_{n=0}^{T(D-1)}a_{n}\,\omega_{m}^{n}\,s^{-n}.$ (22)
@@
$\frac{(1-N^{-1})^{T}}{(1-\omega_{m}s^{-1})^{T}}=\sum_{n=0}^{T(D-1)}a_{n}\,\omega_{m}^{n}\,s^{-n}.$ (23)
@@
$\frac{1}{D}\sum_{m=0}^{D-1}\frac{(1-N^{-1})^{T}}{(1-\omega_{m}s^{-1})^{T}}$ $=\sum_{n=0}^{T(D-1)}a_{n}\,s^{-n}\left(\frac{1}{D}\sum_{m=0}^{D-1}\omega_{m}^{n}\right)$
$=\sum_{k=0}^{S}a_{kD}\,s^{-kD}$
$=\sum_{k=0}^{S}a_{kD}\,N^{-k},$ (24)
@@
Since (6) is a finite sum of nonnegative integer powers of NN with coefficients $a_{kD}\geq 0$ for all $k\in[0:S]$, it follows that $L$ is a polynomial in NN with nonnegative coefficients. Moreover, the highest power of NN occurs at $k=0$, and
$a_{0}=[x^{0}](1+x+\cdots+x^{D-1})^{T}=1,$